# Wi-CBR: WiFi-based Cross-domain Behavior Recognition via Multimodal Collaborative Awareness

### v1

@@ 3.4 Cross-modal Synergistic Interaction and Fusion
We multiply input features $X_{\text{PD}}$ by $W_{1}$ and $W_{2}$ respectively, yielding two weighted features: the strengthen ones $X_{\text{PD}}^{S}$ and less informative ones $X_{\text{PD}}^{W}$.Thus, we completed the interaction between the input phase and DFS features. The attention weights are obtained by learning the variance and bias through group normalization, which is gated to obtain the strengthened and weakened attention maps. Two components obtained are as follows: $X_{\text{PD}}^{S}$ has informative and expressive spatial content and is strengthened, while $X_{\text{PD}}^{W}$ has little or no information, which is considered redundant and weakened.
@@
Cross-modal Fusion. We propose a fusion operation to achieve synergistic utilization. An information-rich feature after strengthening is added to a feature with less information after weakening. New features with richer information are generated, i.e., one feature dominates while the other aids. Instead of adding the two components directly, we use a cross-fertilization operation to thoroughly combine the two weighted different information features to enhance the information flow between them. The cross-rendered features Xw1 and Xw2 are then stitched together to obtain a spatially fine feature mapping of Xw. The whole process of Fusion operation can be expressed as :
$\begin{cases}X_{\text{PD}}^{S}=W_{1}\otimes X_{\text{PD}},X_{\text{PD}}^{W}=W_{2}\otimes X_{\text{PD}};\\ Y_{\text{PS\_DW}}=X_{\text{P}}^{S}\oplus X_{\text{D}}^{W},Y_{\text{PW\_DS}}=X_{\text{P}}^{W}\oplus X_{\text{D}}^{S};\\ X_{\text{out}}=Y_{\text{PS\_DW}}\cup Y_{\text{PW\_DS}},\end{cases}$ (16)
@@
where $\otimes$ is element-wise multiplication, $\oplus$ is element-wise summation, $\cup$ is concatenation. $Y_{\text{PS\_DW}}$ is a phase-dominant DFS-assisted fusion feature, while $Y_{\text{PW\_DS}}$ is a phase-assisted DFS-dominant fusion feature. After the Cross-Model Fusion is applied to the intermediate input features $X_{\text{PD}}$, not only do we separate the informative features from less informative ones, but also we reconstruct them to enhance the representative features and suppress the redundant features in spatial dimension.
3.5 Gestrue Recognition and Contrastive Loss Optimization
@@
$\hat{y}=f_{\text{classifier}}(X_{\text{out}};\Theta_{\text{classifier}}),$ (17)
$\mathcal{L}_{ce}=\mathcal{L}_{ce}(y,\hat{y})=\frac{1}{M}\sum_{m=1}^{M}\sum_{s=1}^{S}y_{m,s}\log(\hat{y}_{m,s}),$ (18)
@@
where $\Theta_{classifier}$ is the learned parameter, $L_{ce}$ is a classification cross-entropy loss, and $M$ is the number of data samples, $S$ is the number of gestures.
Contrastive Loss optimizes the objective by learning a distance measure based on multiple positive and negative sample-to-sample pairs. The key idea behind this is to learn an embedding space where similar pairs of samples are close to each other and dissimilar pairs are far apart. Thus, we can obtain an invariant representation across different environments for the same instance, i.e., a domain-independent characterization. We aim to learn generalizable representations by minimizing the following contrast-based objective.For the $i$th sample extracted feature $x_{i}$:
$\mathcal{L}_{con}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{e^{(x_{i}^{T}x_{+})/\tau}}{e^{(x_{i}^{T}x_{+})/\tau}+\sum e^{(x_{i}^{T}x_{-})/\tau}},$ (19)
@@
where, the pair $(x_{i},x_{+})$ represents samples of the same class (positive pair), while $(x_{i},x_{-})$ denotes samples from different classes (negative pair), with $\tau$ serving as the temperature parameter. Several elements can influence the variation patterns of the CSI signal, with varying degrees of effect. For example, a user’s influence on signal patterns is less significant compared to changes in location and orientation. Importantly, positive sample pairs across diverse environments differ, and some may be challenging to match due to substantial data discrepancies. Perfectly aligning all negative samples might limit the model’s ability to generalize. To mitigate this, we utilize class proxies to symbolize each gesture, ideally enhancing resilience across samples from varied settings. Formally, these proxy vectors are defined as the weights of the final fully connected layer in the classifier. To further improve semantic consistency, we implement a proxy-based contrastive loss that utilizes the connections between class proxies and samples to foster robust representations  [45]. Given the representation $x_{i}$ of $i$th sample , we select its class proxy $w_{c}$ in place of positive samples $x_{+}$ to form proxy-to-sample positive pairs. The contrast loss is incorporated into the overall loss function $\mathcal{L}$:
$\mathcal{L}_{con}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{e^{(w_{c}^{T}xi)/\tau}}{e^{(w_{c}^{T}x_{i})/\tau}+\sum_{k=1,k\neq c}^{R}e^{(w_{k}^{T}x_{i})/\tau}};\quad\mathcal{L}=\mathcal{L}_{ce}+\mathcal{L}_{con},$ (20)
@@
where $w_{c}$ represents the class proxy of sample $i$, $R$ is the number of classes, and $P$ is the number of negative pairs. In general, we aim to minimize the following final loss function, where $\mathcal{L}_{ce}$ is cross-entropy loss, and $\mathcal{L}_{con}$ is the proxy-based contrastive loss. $\beta$ is trade-off parameter.
4 Experiments
4.1 Implementation Details
Datasets Setting. To evaluate the effectiveness of our model for cross-domain gesture recognition, we conducted extensive experiments on the Widar 3.0 and XRF55 datasets. For in-domain, cross-location, and orientation evaluations on Widar 3.0, we used 80% of the data for training and 20% for testing with five-fold cross-validation. For cross-location evaluation, one location was used for testing and the remaining four for training. The in-domain and cross-direction evaluations followed a similar approach. For cross-environment evaluation, we used data from three environments, totaling 12,750 samples (17 users × 5 locations × 5 directions × 6 gestures × 5 instances), with training on two environments and testing on the third using triple cross-validation. For XRF55, the cross-environment evaluation involved four scenarios, with quad-fold cross-validation on 6,240 samples.
In our implementation, we used Matlab to preprocess CSI data and generate 224×224 RGB images. After obtaining phase and DFS images, all models were implemented in PyTorch 1.13.1. The network architecture and data dimensions are shown in Fig. 3. Wi-CBR employs ResNet-18 [46] with pre-trained ImageNet weights as the feature extractor, addressing parameter learning challenges due to limited training data [37, 47]. The Cross-Model Interaction module uses 4 groups, with a threshold set at 0.5. The contrast loss weight $\beta$ and temperature are both set to 0.1. During training, the model is optimized using Adam with a learning rate of 0.001, a batch size of 10, and 30 epochs. We used the same network structure for both the Widar 3.0 and XRF55 datasets, ensuring robustness across datasets. The final fully connected layer is adjusted to 6 or 8 units, corresponding to the number of gestures in each dataset. A random seed of 42 was set for reproducibility.
4.2 Overall Performance
@@
WiSR [50] CSI image 67.73% 69.74% 52.77% 63.41% 26.66%
Recurrent ConFormer [51] Raw CSI 73.84% 85.88% 50.38% 70.03% 16.54%
THAT [52] Raw CSI 71.56% 81.76% 49.71% 67.68% 23.23%
@@
[…]
Comparison with state-of-the-art methods. We first evaluate the overall performance of Wi-CBR in cross-location, cross-orientation, and cross-environment scenarios, compared with some cross-domain recognition methods, as shown in Tab. 1. Modeling-Based method: Widar3.0 [7],WiHF [8],WIGNN [53].They extract DFS features from denoised CSI data, and some further manually extract domain-independent features such as BVP and MCP. The cross-domain average performance of these methods on the Widar3.0 dataset can reach 85.12%. Low Level Semantic Learning-Based method: ImgFi [48],EI [49],WiSR [50],Recurrent ConFormer [51],THAT [52].They use low-level semantic data, such as raw CSI, amplitude, and preliminary CSI visualization images, which can be found. The average cross-domain performance of these efforts on the Widar 3.0 dataset ranges from 40% to 70%, while switching to the XRF55 dataset, the maximum is only up to 31.90%. High Level Semantic Learning-Based method: WIGRUNT [37],WiDual [54],AaD [55], Wi-SFDAGR [38].They use CSI-ratio for the denoising extraction phase and then use an efficient ResNet18 backbone network as a feature extractor. In particular, Wi-SFDAGR uses attraction-dispersion networks to optimize prediction boundaries and enhance feature aggregation. The cross-domain average performance reaches 96.66% on Widar 3.0.
Table 2: Detailed performance on Widar 3.0 and XRF55.
Dataset Setting 1 2 3 4 5
Widar 3.0 In-Domain 99.04% 99.70% 99.48% 99.78% 99.70%
CL 97.85% 97.63% 98.74% 98.00% 99.48%
CO 92.07% 98.22% 97.85% 99.19% 94.15%
CE 93.19% 99.33% 97.20% -- --
XRF55 CE 55.65% 68.13% 72.50% 67.92% --
Wi-CBR generally outperforms state-of-the-art work on both datasets.In addition, we observe an interesting phenomenon: The performance across environments on the XRF55 dataset is lower than that of Widar3.0, regardless of which method is used. This is because the XRF55 dataset has a lower sampling rate than Widar3.0, and the number of receiver arrangements in the environment is half that of Widar3.0. Wi-CBR, on the other hand, demonstrates robustness across datasets, and our work still achieves 66.05% accuracy despite changes in conditions such as receiver and sampling rate.
For the Widar 3.0 dataset, Wi-CBR has an average recognition accuracy of 99.54% for in-domain gesture recognition and 98.34%, 96.30%, and 96.57% for cross-location, orientation, and environment, respectively. For the XRF55 dataset, the average recognition accuracy across environments is 66.05%.The detailed results are shown in Tab. 2.The Setting denotes the data for testing while others for training.
Figure 4: The confusion matrices of our method in the indomain (repetition 1),CL (location 1), CO (orientation 1), and CE (environment 3) settings.
To analyze recognition accuracy and misclassification rates, we present confusion matrices for test location 1, orientation 1, and environment 3 on Widar 3.0, as shown in Fig. 4. In cross-location and cross-orientation tests, “push” and “zigzag” gestures achieve the highest accuracy, while “slide” and “sweep” show the lowest, respectively. In the cross-environment test, “sweep” and “clap” perform best. Interestingly, “slide” is frequently misclassified as “clap.” This likely stems from similar motion trajectories, as both gestures involve inclined or sliding movements, making them harder to distinguish.
Table 3: Ablation Analysis in Different Settings
Methods In-Domain CL CO CE
Wi-CBR 99.54% 98.34% 96.30% 96.57%
Wi-CBR w/o D 99.42% 96.07% 92.76% 93.76%
Wi-CBR w/o P 97.13% 95.04% 87.09% 94.93%
Wi-CBR w/o IF 99.11% 97.07% 95.07% 95.93%
Wi-CBR w/o CL 99.49% 97.26% 95.56% 96.08%
Ablation Study. We evaluate four variants of our model to conduct an ablation study. Wi-CBR w/o D: Only the CSI ratio phase is used as input data for the model. Wi-CBR w/o P: Only the DFS image is utilized as input data. Wi-CBR w/o IF: No interaction, use WiGRUNT  [37] channel-attention module CRU instead of ours. Wi-CBR w/o CL: Only use cross-entropy loss.
The results in Tab. 3 show that input data significantly affects model performance, with phase data and DFS both contributing to recognition accuracy. Phase data provides comprehensive raw information, while DFS has some omissions. Removing our feature interaction and fusion modules consistently reduced model accuracy, highlighting the importance of cross-modal interaction for balancing phase and DFS data. The multimodal fusion module effectively integrates different modal features. Additionally, omitting contrast loss guidance led to a noticeable drop in cross-domain performance, while intra-domain performance remained stable, indicating that contrast loss helps the model focus on gesture-related features rather than environmental noise. Overall, Wi-CBR effectively leverages full CSI raw data and DFS for improved cross-domain performance.
5 Conclusion
We propose Wi-CBR, a WiFi-based cross-domain human behavior recognition framework with strong cross-domain performance. It combines model-based and learning-based approaches, leveraging phase data (dynamic path length changes) and DFS data (frequency changes related to gesture speed). By utilizing full data and domain-independent feature guidance, Wi-CBR enhances recognition accuracy. The framework includes a two-branch self-attention module to capture spatio-temporal cues, followed by bi-gating for feature enhancement and suppression. Cross-modal data fusion in both time and frequency domains enables robust behavioral feature extraction, focusing on the behavior itself while minimizing environmental influence.
Appendix A More Implementation Details
Dataset Details.

### v4

@@ Introduction
We multiply input features $X_{\text{PD}}$ by $G_{1}$ and $G_{2}$ respectively, yielding two weighted features: the strengthen ones $X_{\text{PD}}^{S}$ and less informative ones $X_{\text{PD}}^{W}$.Thus, we completed the interaction between the input phase and DFS features. The attention weights are obtained by learning the variance and bias through group normalization, which is gated to obtain the strengthened and weakened attention maps. Two components obtained are as follows: $X_{\text{PD}}^{S}$ has informative and expressive spatial content and is strengthened, while $X_{\text{PD}}^{W}$ has little or no information, which is considered redundant and weakened.
@@
Saliency Adaptive Fusion. We propose a fusion operation to achieve synergistic utilization. An information-rich feature after strengthening is added to a feature with less information after weakening. New features with richer information are generated, i.e., one feature dominates while the other aids. Instead of adding the two components directly, we use a cross-fertilization operation to thoroughly combine the two weighted different information features to enhance the information flow between them. The cross-rendered features $Y_{\text{PS\_DW}}$ and $Y_{\text{PW\_DS}}$ are then stitched together to obtain a spatially fine feature mapping of $X_{\text{out}}$. The whole process of Fusion operation can be expressed as :
$\begin{cases}X_{\text{PD}}^{S}=G_{1}\otimes X_{\text{PD}},X_{\text{PD}}^{W}=G_{2}\otimes X_{\text{PD}};\\ Y_{\text{DS\_PW}}=X_{\text{D}}^{S}\oplus X_{\text{P}}^{W},Y_{\text{DW\_PS}}=X_{\text{D}}^{W}\oplus X_{\text{P}}^{S};\\ X_{\text{out}}=\mathrm{Concat}(Y_{\text{DS\_PW}},Y_{\text{DW\_PS}})\end{cases}$ (15)
@@
where $\otimes$ is element-wise multiplication, $\oplus$ is element-wise summation, $\mathrm{Concat}$ is concatenation. When DFS is remarkable as a domain-independent feature, we use the enhanced DFS feature, assisted by the weakened detailed phase feature, namely $Y_{\text{DS\_PW}}$. When DFS is weak as a domain-independent feature, we use the weakened DFS feature and use the enhanced detailed phase feature as a supplement, that is, $Y_{\text{DW\_PS}}$. After the Saliency Guidance Module is applied to the input features $X_{\text{PD}}$, not only do we separate the informative features from less informative ones, but also we reconstruct them to enhance the representative features and suppress the redundant features in spatial dimension.
Contrastive Loss Optimization
@@
$\mathcal{L}_{ce}=\mathcal{L}_{ce}(y,\hat{y})=-\frac{1}{M}\sum_{m=1}^{M}\sum_{s=1}^{S}y_{m,s}\log(\hat{y}_{m,s})$ (16)
@@
where $L_{ce}$ is a classification cross-entropy loss, and $M$ is the number of data samples, $S$ is the number of gestures.
@@
Contrastive Loss. optimizes the objective by learning a distance measure based on multiple positive and negative sample-to-sample pairs. The key idea behind this is to learn an embedding space where similar pairs of samples are close to each other and dissimilar pairs are far apart. Thus, we can obtain an invariant representation across different environments for the same instance, i.e., a domain-independent characterization. Several elements can influence the variation patterns of the CSI signal, with varying degrees of effect. For example, a user’s influence on signal patterns is less significant compared to changes in location and orientation. Importantly, positive sample pairs across diverse environments differ, and some may be challenging to match due to substantial data discrepancies. Perfectly aligning all samples might limit the model’s ability to generalize. To mitigate this, we utilize class proxies to symbolize each gesture, ideally enhancing resilience across samples from varied settings. Formally, these proxy vectors are defined as the weights of the final fully connected layer in the classifier. To further improve semantic consistency, we implement a proxy-based contrastive loss that utilizes the connections between class proxies and samples to foster robust representations  (Proxy). Given the representation $x_{i}$ of the $i$th sample, we select its class proxy $w_{c}$ in place of positive samples $x_{+}$ to form proxy-to-sample positive pairs. The contrast loss is incorporated into the overall loss function $\mathcal{L}$:
$\mathcal{L}_{con}$ $=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{e^{(w_{c}^{T}xi)/\tau}}{e^{(w_{c}^{T}x_{i})/\tau}+\sum_{k=1,k\neq c}^{R}e^{(w_{k}^{T}x_{i})/\tau}};$
$\mathcal{L}=\mathcal{L}_{ce}+\beta\times\mathcal{L}_{con}$ (17)
@@
where $w_{c}$ represents the class proxy corresponding to class $c$, $R$ is the total number of classes, and $\tau$ serves as the temperature parameter. We aim to minimize the following final loss function, where $\mathcal{L}_{ce}$ is the cross-entropy loss and $\mathcal{L}_{con}$ is the proxy-based contrastive loss. $\beta$ is the trade-off parameter.
@@
THAT (THAT) Raw CSI 71.56 81.76 49.71 67.68 23.23
[…]
WiSR (WISR) CSI image 67.73 69.74 52.77 63.41 26.66
Recurrent ConFormer (Recurrent_ConFormer) Raw CSI 73.84 85.88 50.38 70.03 16.54
[…]
Experiments
Experimental Setup
We train and evaluate our proposed Wi-CBR on the two largest human sensing multimodal public datasets.
Gesture Recognition. To evaluate the effectiveness of our model for cross-domain gesture recognition, we conducted extensive experiments on the Widar 3.0 and XRF55 datasets. For in-domain, cross-location, and orientation evaluations on Widar 3.0, we used 80% of the data for training and 20% for testing with five-fold cross-validation. For cross-location evaluation, one location was used for testing and the remaining four for training. The in-domain and cross-direction evaluations followed a similar approach. For cross-environment evaluation, we used data from three environments, totaling 12,750 samples, with training on two environments and testing on the third using triple cross-validation. For XRF55, the cross-environment evaluation involved four scenarios, with quad-fold cross-validation on 6,240 samples.
Activity Recognition. For the fairness of the experimental evaluation, the implementation details are exactly the same as those of XRF55. Note that the quad-fold cross-validation is not used here. At this time, due to the lack of multi-environment data support, the cross-domain migration capability of the model is highly required. The training set of scene 1 is used for training, that is, the first 14 times of each behaviour of each user. The test set is tested with the samples of scenes 2, 3, and 4, respectively. There are 33000 samples in the training set and 3300 samples in each scene in the test set, including 55 daily human behaviours.
Implementation Details. We used MATLAB to preprocess CSI data and generate 224×224 RGB images. After obtaining phase and DFS images, all models were implemented in PyTorch 1.13.1. The network architecture and data dimensions are shown in Fig. 2. Wi-CBR employs ResNet-18 (60) with pre-trained ImageNet weights as the feature extractor. The Cross-Model Interaction module set the threshold to 0.5. The contrast loss weight $\beta$ and temperature are both set to 0.1. During training, the model is optimized using Adam with a learning rate of 0.0001, a batch size of 10, and 30 epochs. We used the same network structure for both the Widar 3.0 and XRF55 datasets, ensuring robustness across datasets. Full connection layer classification header: in gesture recognition, it is 6 to 9 in widar3.0 dataset, 8 in XRF55 dataset, and 55 in activity recognition. A random seed of 42 was set for reproducibility.
Comparisons to Prior SOTA Results
We compare our method with the baseline method in downstream human perception tasks, including fine-grained HGR and human HAR. We mainly study single-factor cross-domain, that is, only one domain factor we have never seen in location, orientation, and environment. We trained in some specific domains and test in an unknown domain.
CE Method few-shot
zero one two
2 WIGRUNT(WIGRUNT) 21.82 41.34 50.17
XRF55(XRF55) 2.52 49.83 57.51
Wi-CBR (Ours) 31.33 50.92 58.96
3 WIGRUNT(WIGRUNT) 13.12 41.34 56.53
XRF55(XRF55) 2.14 50.85 63.30
Wi-CBR (Ours) 28.60 51.74 61.01
4 WIGRUNT(WIGRUNT) 14.55 42.60 53.81
XRF55(XRF55) 2.03 50.94 61.41
Wi-CBR (Ours) 31.30 52.70 62.66
Table 2: 55 kinds of human daily behaviors cross-domain recognition on XRF55 under few-shot.
Figure 3: The confusion matrices in the ID (repetition 1), CL (location 1), CO (orientation 1), and CE (environment 3).
HGR Comparison with SOTA Methods. We compare the cross-domain performance with SOTA methods across modeling-based, low-level semantic, and high-level semantic approaches (Tab. 1). Modeling-based methods (e.g., Widar3.0 (Widar3.0), WiHF (WIHF)) extract DFS and domain-independent features, achieving 85.12% on Widar3.0. Low-level semantic methods (e.g., ImgFi (Imgfi), WiSR (WISR)) rely on raw CSI or visualizations, with performance ranging from 40% to 70% on Widar3.0 and up to 31.90% on XRF55. High-level semantic methods (e.g., Wi-SFDAGR (Wi-SFDAGR)) utilize CSI-ratio denoising and advanced networks like ResNet18, reaching 96.66% on Widar3.0 through optimized feature aggregation.
Wi-CBR outperforms state-of-the-art on both datasets. In addition, we observe an interesting phenomenon: The performance across environments on the XRF55 dataset is lower than that of Widar3.0, regardless of which method is used. This is because the XRF55 dataset has a lower sampling rate than Widar3.0, and the number of receiver arrangements in the environment is half that of Widar3.0. Wi-CBR, on the other hand, demonstrates robustness across datasets, and our work still achieves 66.05% accuracy despite changes in conditions such as receiver and sampling rate.
Methods ID CL CO CE
Without DFS 99.42 96.07 92.76 93.76
Without Phase 97.13 95.04 87.09 94.93
Without CL 99.49 97.26 95.56 96.08
Table 3: Effectiveness of each component.
HAR included Comparison on XRF55. We further evaluate the recognition of human daily activities. We initially train the model using training samples from environment 1, and subsequently finetune it using 0/1/2 samples (per subject per action) from environment 2. We then test the finetuned model with the remaining 20/19/18 samples (per subject per action) in environment 2. Tab. 2 indicates the model’s best transferability to unseen environment. In the case of zero-shot, most models can not effectively learn the domain invariance feature, and the model is close to a random guess. The advantage of Wi-CBR is that it can recognize the good features learned by multimodal collaborative sensing even if it is only learned in one environment, without using any samples of an unknown environment.
Ablation Study
We evaluated the impact of different components and fusion methods on the experiment. The results in Tab. 3 show that DFS greatly improves cross-environment performance, while phase mainly affects in-domain and cross-domain performance within the environment. Additionally, omitting contrast loss guidance led to a noticeable drop in cross-domain performance, while in-domain performance remained stable, indicating that contrast loss helps the model focus on gesture-related features rather than environmental noise. The results in Tab. 4 show that easy channel attention using WiGRUNT  (WIGRUNT) channel-gate module will continue to reduce the accuracy of the model, highlighting the importance of interaction for balancing phase and DFS data. We also tried different fusion methods. In fact, when there is only one kind of data as input, it is equivalent to self-interaction, such as using PS_PW. When using the two kinds of data as input, the respective enhanced addition PS_DS exists feature redundancy and cannot play a complementary effect. Overall, Wi-CBR effectively leverages full CSI raw data and DFS for improved cross-domain performance.
Methods ID CL CO CE
Channel Attention 99.11 97.07 95.07 95.93
PS_DS and PW_DW 99.54 97.14 95.13 96.28
Wi-CBR 99.54 98.34 96.30 96.57
Table 4: Effectiveness of Fusion.
To analyze recognition accuracy and misclassification rates, we present confusion matrices for test location 1, orientation 1, and environment 3 on Widar 3.0, as shown in Fig. 3. In cross-location and cross-orientation tests, “push” and “zigzag” achieve the best accuracy, while “slide” and “sweep” show the lowest, respectively. In the cross-environment test, “sweep” and “clap” perform best. Interestingly, “slide” is frequently misclassified as “clap.” This likely stems from similar motion trajectories, as both gestures involve inclined or sliding movements, making them harder to distinguish.
Conclusion
In this work, we propose Wi-CBR, a WiFi-based cross-domain human behavior recognition framework with strong cross-domain performance. It integrates model-based and learning-based methods, using phase data (path length changes) and DFS data (frequency shifts) for robust cross-domain performance. It employs a two-branch self-attention module to extract spatio-temporal features and a saliency-aware adaptive mechanism to enhance phase features based on DFS significance. This time-frequency fusion focuses on behavior while reducing environmental interference, improving recognition accuracy.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (Grants No. 62502144, 61932009, 62462015), the Anhui Provincial Natural Science Foundation, China (Grant No. 2408085QF191), the Young Elite Scientist Sponsorship Program by Gast (Grant No. GASTYESS202429), and the Fundamental Research Funds for the Central Universities (Grants No. JZ2024HGTA0178, PA2025IISL0112). The computation is completed on the HPC Platform of Hefei University of Technology.
Appendix
Appendix A A. More Experimental Details
A.1 More Experiment Results
For the Widar 3.0 dataset, Wi-CBR has an average recognition accuracy of 99.54% for in-domain gesture recognition and 98.34%, 96.30%, and 96.87% for cross location, orientation, and environment, respectively. For the XRF55 dataset, the average recognition accuracy across environments is 66.05%. The detailed results are shown in Tab. 5. The Setting denotes the data for testing, while others are for training. We report the results on accuracy and macro F1 with additional random seeds as Tab. 6.
Dataset Setting 1 2 3 4 5
Widar 3.0 In-Domain 99.04 99.70 99.48 99.78 99.70
CL 97.85 97.63 98.74 98.00 99.48
CO 92.07 98.22 97.85 99.19 94.15
CE 93.04 99.40 98.16 - -
XRF55 CE 55.65 68.13 72.50 67.92 -
Table 5: Detailed gesture recognition performance on Widar 3.0 and XRF55.
@@ A.1 More Experiment Results
Table 6: Results under different random seeds.
A.2 Impact of number of users and gestures
We evaluate Wi-CBR under varying numbers of users (default: 9) and gestures (default: 6), with results in Tab. 7 and Tab. 8, respectively. As users increase, Wi-CBR generally improves , indicating better adaptation to more users, whereas WiGRUNT declines, with a sharp drop when adding the first extra user; overall, Wi-CBR benefits from more users. As gestures increase, both methods degrade; however, Wi-CBR drops more slowly than WiGRUNT for in-domain and cross-location and can rebound on easily recognized gestures, while both show similar trends under cross-orientation. WiGRUNT decreases monotonically, suggesting Wi-CBR is more stable as gesture classes grow.
No. of Users 6 7 8 9
WiGRUNT In-Domain 99.60 99.45 99.48 99.10
CL 96.49 96.21 96.72 96.17
CO 93.36 92.86 94.11 93.95
Wi-CBR In-Domain 99.07 99.54 99.63 99.54
CL 97.20 98.46 98.92 98.34
CO 95.31 96.84 97.08 96.57
Table 7: Accuracy of WiGRUNT and Wi-CBR for different numbers of users.
No. of Gestures 6 7 8 9
WiGRUNT In-Domain 99.60 99.26 98.32 98.08
CL 96.49 95.45 92.25 93.58
CO 93.36 92.12 89.22 89.01
Wi-CBR In-Domain 99.07 99.31 98.67 98.61
CL 97.20 97.52 95.77 96.03
CO 95.31 94.89 91.20 91.11
Table 8: Accuracy of WiGRUNT and Wi-CBR for different numbers of gestures.
Appendix B B. More Dataset Details
B.1 Gesture Recognition Dataset Details